# Manta: Enhancing Mamba for Few-Shot Action Recognition of Long Sub-Sequence

### v1

@@ Abstract
In few-shot action recognition (FSAR), long sub-sequences of video naturally express entire actions more effectively. However, the computational complexity of mainstream Transformer-based methods limits their application. Recent Mamba demonstrates efficiency in modeling long sequences, but directly applying Mamba to FSAR overlooks the importance of local feature modeling and alignment. Moreover, long sub-sequences within the same class accumulate intra-class variance, which adversely impacts FSAR performance. To solve these challenges, we propose a Matryoshka MAmba and CoNtrasTive LeArning framework (Manta). Firstly, the Matryoshka Mamba introduces multiple Inner Modules to enhance local feature representation, rather than directly modeling global features. An Outer Module captures dependencies of timeline between these local features for implicit temporal alignment. Secondly, a hybrid contrastive learning paradigm, combining both supervised and unsupervised methods, is designed to mitigate the negative effects of intra-class variance accumulation. The Matryoshka Mamba and the hybrid contrastive learning paradigm operate in parallel branches within Manta, enhancing Mamba for FSAR of long sub-sequence. Manta achieves new state-of-the-art performance on prominent benchmarks, including SSv2, Kinetics, UCF101, and HMDB51. Extensive empirical studies prove that Manta significantly improves FSAR of long sub-sequence from multiple perspectives. The code is released at https://github.com/wenbohuang1002/Manta.
@@ Appendix A Supplementary Materials
Due to the higher accuracy and better efficiency, we select the up-to-date Mamba-2 (Dao and Gu 2024) as the base model of our Inner Module and Outer Module. To compare the performance under different base models, we also conducted experiments with Mamba-1 (Gu and Dao 2023). From the results demonstrated in Table 6, we observe that Mamba-2 performs better than Mamba-1. The reason is that Mamba-2 enlarges the state space to 16 or even larger and integrates the attention mechanism (Vaswani et al. 2017), resulting in better performance Therefore, we identify Mamba-2 as the base model.
@@
Table 6: Comparison ($\uparrow$ Acc. %) between Mamba-1 (M-1) and Mamba-2 (M-2). Few-shot setting: 5-way 1-shot.
@@
Inner Module is applied for reinforcing local features while Outer Module is designed for executing implicit temporal alignment. Exploring independent and shared parameters between bub-branches can improve the differentiation of their functions. Therefore, we complete experiments by various forms of the parameters and list the results in Table 7. Using all independent or shared parameters does not provide the best performance. Inner Module with independent parameters and Outer Module with shared parameters can achieve the best performance. We suppose that the variable local features need more flexible bidirectional scanning from Inner Module with independent parameters. While temporal alignment is a bidirectional symmetric structure, Outer Module with shared parameters can make alignment learned in both directions reinforce each other.
@@
Table 7: Comparison ($\uparrow$ Acc. %) between independent (I-P) and shared parameters (S-P). Few-shot setting: 5-way 1-shot.
@@
To discuss the difference between various methods of multi-scale combination, we compare fixed and learnable weights in our experiments. From the results revealed in Table 8, we find that applying learnable weights ($w_{o}$) performs better than directly using fixed ones ($w_{o}$). The reason is that weights allow the model to dynamically adjust appropriate significance to the features at each scale, thus enhancing the overall feature representation.
@@
Table 8: Comparison ($\uparrow$ Acc. %) between fixed and learnable weights $w_{o}$. Few-shot setting: 5-way 1-shot.
@@
We intuitively believe that temporal order is crucial for feature representation in FSAR. To verify this conjecture, we reverse the order of the support set while maintaining the temporal order of the query, and then conduct experiments for comparison. The results demonstrated in Table 9 show that all datasets have a degeneration with reserved order. It is worth noting that the drop in SSv2 is more significant than Kinetics. The dataset type contributes to our observation that temporal-related SSv2 is more sensitive to changes in temporal order. Therefore, the negative impact of reserved order is smaller on Kinetics.
@@
Table 9: Comparison ($\uparrow$ Acc. %) between reserved and temporal order. Few-shot setting: 5-way 1-shot.
@@
The weight factor $\lambda$ of loss is designed to adjust values of $\mathcal{L}_{\text{ce}}$ to the same level of $\mathcal{L}_{\text{hc}}$. We aim to find the most suitable $\lambda$ from extensive experiments. The results are recorded in Table 10. With the increase of $\lambda$, the performance gradually improves. However, we observe that the best results always be achieved when $\lambda=4$. If the value of $\lambda$ continues to grow, there will be a small decline in performance. Therefore, we determined that $\lambda=4$ is the most appropriate weight factor.
@@
Table 10: Comparison ($\uparrow$ Acc. %) of various weight factor $\lambda$.
@@
After obtaining $\mathcal{L}_{\text{hc}}$ from $\mathcal{L}_{\text{con}}^{\mathcal{S}}$, $\mathcal{L}_{\text{con}}^{\mathcal{Q}}$, and $\mathcal{L}_{\text{con}}^{\mathcal{SQ}}$, we aim to explore the effect of each component. By analyzing $\mathcal{L}_{\text{hc}}$ in detail, we conducted experiments to evaluate the impact of each contrastive loss term. Results presented in Table 11 reveal that using any single type of contrastive learning is beneficial for performance improvement. However, employing both contrastive learning losses together yields better results than using just one. Finally, applying the full $\mathcal{L}_{\text{hc}}$ in a hybrid manner achieves the best performance. This observation confirms that all components of $\mathcal{L}_{\text{hc}}$ are necessary.
@@
Table 11: Comparison ($\uparrow$ Acc. %) of various contrastive loss.
@@
After determining the hybrid way of $\mathcal{L}_{\text{hc}}$, the impact of hyper-parameter temperature $\tau$ is also discussed. By adjusting $\tau$ from 0.01 to 0.11, we conducted several experiments and reported results in Table 12. The overall trend of performance is gradually improved with the $\tau$ growth. However, the improvement is not unbounded, peaking at $\tau=0.07$ and saturating. Further growth of $\tau$ makes the performance drop. Therefore, the temperature $\tau$ is determined to be 0.07.
@@
Table 12: Comparison ($\uparrow$ Acc. %) of various temperature $\tau$.
@@
From the testing set of Kinetics, 5 challenging action classes with similar appearances including “Unboxing”, “Folding paper”, “Hurling (sport)”, “Tap dancing”, and “Throwing axe” is selected. In each class, there are several samples with significant intra-class variance. From the t-SNE (Van der Maaten and Hinton 2008), feature representation without Manta (Figure 8(a)) has overlap between classes because of the similar visual features of actions. Due to the negative impact of intra-class variance accumulation, many samples are separated and cannot be clustered well. The above phenomena can be alleviated with the help of Manta (Figure 8(b)). Specifically, the different classes are separated and the samples belonging to the same class are tightly clustered. For these independently separated samples with significant intra-class variances, a tendency of clustering to their corresponding class is highly remarkable.
@@
Figure 8: Challenging action classes are selected from Kinetics. Blue: “Unboxing”, Orange: “Folding paper”, Green: “Hurling (sport)”, Red: “Tap dancing”, Purple: “Throwing axe”. In the t-SNE visualization, points with the same color are samples of the same class, and points with black borders are samples with large intra-class variance. Few-shot setting: 5-way 25-shot.
@@
Although long sub-sequence has inherent advantage in expressing entire actions, directly applying Mamba for FSAR of long sub-sequence cannot obtain satisfactory performance We select 250 samples from 5 classes of UCF101 and visualize the effect of Manta by confusion matrix. As illustrated in Figure 9, “Basketball” and “Basketball dunk” can be seen as actions cover “High jump”. “Soccer juggling”, and “Soccer penalty” can be treated as similar composite actions. We observe Manta reduces the misclassification between them to a lower level. The accuracy of these classes are improved from 55.6% to 89.2%. The experimental results demonstrates the superior of Manta for FSAR of long sub-sequence.
@@
Figure 9: Confusion Matrices of 250 samples from 5 action classes. Ground truth $y_{Q}^{j}$: $X$-axis, predict labels $\tilde{y}_{Q}^{j}$: $Y$-axis. From top to bottom and left to right, action class are “High jump”, “Basketball”, “Basketball dunk”, “Soccer juggling”, and “Soccer penalty” from UCF101. Misclassification be reduced to a low level are highlighted.
@@
During the collection of a large number of samples in the real environment, sample-level noise inevitably occurs for different reasons. Sometimes samples from other classes might be mixed into a particular class. It is time-consuming and laborious to correct them. Therefore evaluating with frame-level noise can simulate more authentic scenarios, testing the robustness of models. As shown in Table 13, we find that the performance of TRX (Perrett et al. 2021), HyRSM (Wang et al. 2022), and MoLo (Wang et al. 2023) decrease severely with the sample-level noise ratio increases. Our Manta is less negatively affected and performs smoothly under sample-level noise. This experiment proves that emphasizing local features, executing temporal alignment, and relieving intra-class variance accumulation can improve robustness.
@@
Table 13: Comparison ($\uparrow$ Acc. %) with sample-level noise. Few-shot setting: 5-way 10-shot.
@@
The real shooting condition not only imposes frame-level noise into samples but also introduces different types of background noise. To be specific, Gaussian noise is related to issues of hardware. Rainy noise is caused by changeable weather. Changes in lighting conditions bring light noise. These all challenge the robustness of the model. Therefore, we added Gaussian, rainy, and light noise to 25% samples respectively to create more intra-class variances, as depicted in Figure 10. The results are listed in Table 14. Consistent with our intuition, Gaussian, rainy, and light noise all negatively affect the performance of the model. Taking the most significant influence from light noise as an example, we observe that our Manta experiences an about 5% performance decline. Compared with other recent methods, Manta performs much more robustly. A similar phenomenon can also be observed in other types of background noise.
Figure 10: Example “Paragliding” from Kinetics for background noise demonstration. Original frames (O) are located in the first row. Gaussian noise (G), rainy noise (R), and light noise (L) are added in the second, third, and fourth rows.
@@
Table 14: Comparison ($\uparrow$ Acc. %) with original frames (O), Gaussian noise (G), rainy noise (R), and light noise (L). Few-shot setting: 5-way 1-shot.
@@
Due to different data distributions, training and testing on various datasets can simulate the performance of models in practical applications. Therefore, we apply Kinetics and SSv2 for cross-dataset testing after eliminating overlapping classes in the training, validation, and test sets. From the results demonstrated in Table 15, our Manta with cross-dataset testing is still ahead of other methods, although this particular test setup can degrade performance. This trend is consistent with regular testing. Compared with other methods, we further find that the performance decline caused by cross-dataset testing is not evident on Manta, proving the better robustness with different data distributions.
@@
Inference with a single GPU can help evaluate the model in practical application with limited resources. To realize such hardware conditions, we select only one 24GB NVIDIA GeForce RTX 3090 GPU in a server for 10,000 tasks in this experiment. As results demonstrated in Table 16, the inference speed of Manta is much faster than other methods based on Transformer. The lower latency of Manta in practice is due to Mamba not only having a more lightweight architecture but also using parallel algorithms optimized for modern hardware, especially GPUs.
@@
Table 16: Inference speed ($\downarrow$ hour) with 10,000 random tasks on single 24GB NVIDIA GeForce RTX 3090 GPU. Few-shot setting: 5-way 10-shot.
@@
For a better comprehension of our Manta, we list the major tensor changes from the Mamba branch in Table 17. The notations are consistent with the main paper. These can demonstrate how Manta works, helping select appropriate hyper-parameters and model settings. According to abundant experiments conducted in previous sections, we determine hyper-parameters such as multi-scale ($\mathcal{O}=\left\{1,2,4\right\}$), temperature ($\tau=0.07$) and weight factor of loss ($\lambda=4$). Model settings including $\mathrm{IM}\left(\cdot\right)$ with independent parameters, $\mathrm{OM}\left(\cdot\right)$ with shared parameters, applying learnable weights $w_{o}$, and hybrid contrastive learning loss $\mathcal{L}_{\text{hc}}$ are more clear. The following pseudo-code provides a better analysis of computational complexity.
@@
Table 17: Model structure of Manta. Major tensor changes are also provided.

### v4

# Long Sub-Sequence

@@ Abstract
In few-shot action recognition (FSAR), long sub-sequences of video naturally express entire actions more effectively. However, the high computational complexity of mainstream Transformer-based methods limits their application. Recent Mamba demonstrates efficiency in modeling long sequences, but directly applying Mamba to FSAR overlooks the importance of local feature modeling and alignment. Moreover, long sub-sequences within the same class accumulate intra-class variance, which adversely impacts FSAR performance. To solve these challenges, we propose a Matryoshka MAmba and CoNtrasTive LeArning framework (Manta). Firstly, the Matryoshka Mamba introduces multiple Inner Modules to enhance local feature representation, rather than directly modeling global features. An Outer Module captures dependencies of timeline between these local features for implicit temporal alignment. Secondly, a hybrid contrastive learning paradigm, combining both supervised and unsupervised methods, is designed to mitigate the negative effects of intra-class variance accumulation. The Matryoshka Mamba and the hybrid contrastive learning paradigm operate in two parallel branches within Manta, enhancing Mamba for FSAR of long sub-sequence. Manta achieves new state-of-the-art performance on prominent benchmarks, including SSv2, Kinetics, UCF101, and HMDB51. Extensive empirical studies prove that Manta significantly improves FSAR of long sub-sequence from multiple perspectives.
@@ Appendix A Supplementary Materials
Due to the higher accuracy and better efficiency, we select the up-to-date Mamba-2 (Dao and Gu 2024) as the base model of our Inner Module and Outer Module. To compare the performance under different base models, we also conducted experiments with Mamba-1 (Gu and Dao 2023). From the results demonstrated in Table I, we observe that Mamba-2 performs better than Mamba-1. The reason is that Mamba-2 enlarges the state space to 16 or even larger and integrates the attention mechanism (Vaswani et al. 2017), resulting in better performance Therefore, we identify Mamba-2 as the base model.
@@
Table I: Comparison ($\uparrow$ Acc. %) between Mamba-1 (M-1) and Mamba-2 (M-2). Few-shot setting: 5-way 1-shot.
@@
Inner Module is applied for reinforcing local features while Outer Module is designed for executing implicit temporal alignment. Exploring independent and shared parameters between bub-branches can improve the differentiation of their functions. Therefore, we complete experiments by various forms of the parameters and list the results in Table II. Using all independent or shared parameters does not provide the best performance. Inner Module with independent parameters and Outer Module with shared parameters can achieve the best performance. We suppose that the variable local features need more flexible bidirectional scanning from Inner Module with independent parameters. While temporal alignment is a bidirectional symmetric structure, Outer Module with shared parameters can make alignment learned in both directions reinforce each other.
@@
Table II: Comparison ($\uparrow$ Acc. %) between independent (I-P) and shared parameters (S-P). Few-shot setting: 5-way 1-shot.
@@
To discuss the difference between various methods of multi-scale combination, we compare fixed and learnable weights in our experiments. From the results revealed in Table III, we find that applying learnable weights ($w_{o}$) performs better than directly using fixed ones ($w_{o}$). The reason is that weights allow the model to dynamically adjust appropriate significance to the features at each scale, thus enhancing the overall feature representation.
@@
Table III: Comparison ($\uparrow$ Acc. %) between fixed and learnable weights $w_{o}$. Few-shot setting: 5-way 1-shot.
@@
We intuitively believe that temporal order is crucial for feature representation in FSAR. To verify this conjecture, we reverse the order of the support set while maintaining the temporal order of the query, and then conduct experiments for comparison. The results demonstrated in Table IV show that all datasets have a degeneration with reserved order. It is worth noting that the drop in SSv2 is more significant than Kinetics. The dataset type contributes to our observation that temporal-related SSv2 is more sensitive to changes in temporal order. Therefore, the negative impact of reserved order is smaller on Kinetics.
@@
Table IV: Comparison ($\uparrow$ Acc. %) between reserved and temporal order. Few-shot setting: 5-way 1-shot.
@@
The weight factor $\lambda$ of loss is designed to adjust values of $\mathcal{L}_{\text{ce}}$ to the same level of $\mathcal{L}_{\text{hc}}$. We aim to find the most suitable $\lambda$ from extensive experiments. The results are recorded in Table V. With the increase of $\lambda$, the performance gradually improves. However, we observe that the best results always be achieved when $\lambda=4$. If the value of $\lambda$ continues to grow, there will be a small decline in performance. Therefore, we determined that $\lambda=4$ is the most appropriate weight factor.
@@
Table V: Comparison ($\uparrow$ Acc. %) of various weight factor $\lambda$.
@@
After obtaining $\mathcal{L}_{\text{hc}}$ from $\mathcal{L}_{\text{con}}^{\mathcal{S}}$, $\mathcal{L}_{\text{con}}^{\mathcal{Q}}$, and $\mathcal{L}_{\text{con}}^{\mathcal{SQ}}$, we aim to explore the effect of each component. By analyzing $\mathcal{L}_{\text{hc}}$ in detail, we conducted experiments to evaluate the impact of each contrastive loss term. Results presented in Table VI reveal that using any single type of contrastive learning is beneficial for performance improvement. However, employing both contrastive learning losses together yields better results than using just one. Finally, applying the full $\mathcal{L}_{\text{hc}}$ in a hybrid manner achieves the best performance. This observation confirms that all components of $\mathcal{L}_{\text{hc}}$ are necessary.
@@
Table VI: Comparison ($\uparrow$ Acc. %) of various contrastive loss.
@@
After determining the hybrid way of $\mathcal{L}_{\text{hc}}$, the impact of hyper-parameter temperature $\tau$ is also discussed. By adjusting $\tau$ from 0.01 to 0.11, we conducted several experiments and reported results in Table VII. The overall trend of performance is gradually improved with the $\tau$ growth. However, the improvement is not unbounded, peaking at $\tau=0.07$ and saturating. Further growth of $\tau$ makes the performance drop. Therefore, the temperature $\tau$ is determined to be 0.07.
@@
Table VII: Comparison ($\uparrow$ Acc. %) of various temperature $\tau$.
@@
From the testing set of Kinetics, 5 challenging action classes with similar appearances including “Unboxing”, “Folding paper”, “Hurling (sport)”, “Tap dancing”, and “Throwing axe” is selected. In each class, there are several samples with significant intra-class variance. From the t-SNE (Van der Maaten and Hinton 2008), feature representation without Manta (Figure I(a)) has overlap between classes because of the similar visual features of actions. Due to the negative impact of intra-class variance accumulation, many samples are separated and cannot be clustered well. The above phenomena can be alleviated with the help of Manta (Figure I(b)). Specifically, the different classes are separated and the samples belonging to the same class are tightly clustered. For these independently separated samples with significant intra-class variances, a tendency of clustering to their corresponding class is highly remarkable.
@@
Figure I: Challenging action classes are selected from Kinetics. Blue: “Unboxing”, Orange: “Folding paper”, Green: “Hurling (sport)”, Red: “Tap dancing”, Purple: “Throwing axe”. In the t-SNE visualization, points with the same color are samples of the same class, and points with black borders are samples with large intra-class variance. Few-shot setting: 5-way 25-shot.
@@
Although long sub-sequence has inherent advantage in expressing entire actions, directly applying Mamba for FSAR of long sub-sequence cannot obtain satisfactory performance We select 250 samples from 5 classes of UCF101 and visualize the effect of Manta by confusion matrix. As illustrated in Figure II, “Basketball” and “Basketball dunk” can be seen as actions cover “High jump”. “Soccer juggling”, and “Soccer penalty” can be treated as similar composite actions. We observe Manta reduces the misclassification between them to a lower level. The accuracy of these classes are improved from 55.6% to 89.2%. The experimental results demonstrates the superior of Manta for FSAR of long sub-sequence.
@@
Figure II: Confusion Matrices of 250 samples from 5 action classes. Ground truth $y_{Q}^{j}$: $X$-axis, predict labels $\tilde{y}_{Q}^{j}$: $Y$-axis. From top to bottom and left to right, action class are “High jump”, “Basketball”, “Basketball dunk”, “Soccer juggling”, and “Soccer penalty” from UCF101. Misclassification be reduced to a low level are highlighted.
@@
During the collection of a large number of samples in the real environment, sample-level noise inevitably occurs for different reasons. Sometimes samples from other classes might be mixed into a particular class. It is time-consuming and laborious to correct them. Therefore evaluating with frame-level noise can simulate more authentic scenarios, testing the robustness of models. As shown in Table VIII, we find that the performance of TRX (Perrett et al. 2021), HyRSM (Wang et al. 2022), and MoLo (Wang et al. 2023) decrease severely with the sample-level noise ratio increases. Our Manta is less negatively affected and performs smoothly under sample-level noise. This experiment proves that emphasizing local features, executing temporal alignment, and relieving intra-class variance accumulation can improve robustness.
@@
Table VIII: Comparison ($\uparrow$ Acc. %) with sample-level noise. Few-shot setting: 5-way 10-shot.
@@
The real shooting condition not only imposes frame-level noise into samples but also introduces different types of background noise. To be specific, Gaussian noise is related to issues of hardware. Rainy noise is caused by changeable weather. Changes in lighting conditions bring light noise. These all challenge the robustness of the model. Therefore, we add Gaussian, rainy, and light noise to 25% samples respectively to create more intra-class variances, as depicted in Figure III. The results are listed in Table IX. Consistent with our intuition, Gaussian, rainy, and light noise all negatively affect the performance of the model. Taking the most significant influence from light noise as an example, we observe that our Manta experiences an about 5% performance decline. Compared with other recent methods, Manta performs much more robustly. A similar phenomenon can also be observed in other types of background noise.
Figure III: Example “Paragliding” from Kinetics for background noise demonstration. Original frames (O) are located in the first row. Gaussian noise (G), rainy noise (R), and light noise (L) are added in the second, third, and fourth rows.
@@
Table IX: Comparison ($\uparrow$ Acc. %) with original frames (O), Gaussian noise (G), rainy noise (R), and light noise (L). Few-shot setting: 5-way 1-shot.
@@
Due to different data distributions, training and testing on various datasets can simulate the performance of models in practical applications. Therefore, we apply Kinetics and SSv2 for cross-dataset testing after eliminating overlapping classes in the training, validation, and test sets. From the results demonstrated in Table X, our Manta with cross-dataset testing is still ahead of other methods, although this particular test setup can degrade performance. This trend is consistent with regular testing. Compared with other methods, we further find that the performance decline caused by cross-dataset testing is not evident on Manta, proving the better robustness with different data distributions.
@@
Inference with a single GPU can help evaluate the model in practical application with limited resources. To realize such hardware conditions, we select only one 24GB NVIDIA GeForce RTX 3090 GPU in a server for 10,000 tasks in this experiment. As results demonstrated in Table XI, the inference speed of Manta is much faster than other methods based on Transformer. The lower latency of Manta in practice is due to Mamba not only having a more lightweight architecture but also using parallel algorithms optimized for modern hardware, especially GPUs.
@@
Table XI: Inference speed ($\downarrow$ hour) with 10,000 random tasks on single 24GB NVIDIA GeForce RTX 3090 GPU. Few-shot setting: 5-way 10-shot.
@@
For a better comprehension of our Manta, we list the major tensor changes from the Mamba branch in Table XII. The notations are consistent with the main paper. These can demonstrate how Manta works, helping select appropriate hyper-parameters and model settings. According to abundant experiments conducted in previous sections, we determine hyper-parameters such as multi-scale ($\mathcal{O}=\left\{1,2,4\right\}$), temperature ($\tau=0.07$) and weight factor of loss ($\lambda=4$). Model settings including $\mathrm{IM}\left(\cdot\right)$ with independent parameters, $\mathrm{OM}\left(\cdot\right)$ with shared parameters, applying learnable weights $w_{o}$, and hybrid contrastive learning loss $\mathcal{L}_{\text{hc}}$ are more clear. The following pseudo-code provides a better analysis of computational complexity.
@@
Table XII: Model structure of Manta. Major tensor changes are also provided.